# How to build the best medical image segmentation algorithm using foundation models: a comprehensive empirical study with Segment Anything Model

### v2

@@ 2.2.4 Summary
To sum up, we choose the following combinations to fine-tune SAM’s performance for specific datasets or tasks, as illustrated in Figure 2: (1) Variations in the image encoder’s architecture, including ViT-H, ViT-B, and ViT-T from MobileSAM [94]. (2) Different approaches to the fine-tuning scope of the network, encompassing both the encoder and decoder updates, as well as the decoder only. (3) Diverse parameter updating settings, comprising Vanilla, PEFT with Adapter blocks, and PEFT with LoRA blocks. By integrating these three variables, we present a total of distinct architectures for fine-tuning SAM on task-specific supervised training setup.
@@
| MRI-BrainSclerosis (MRI-Scler.) | Brain Sclerosis with Lesion Segmentation[63] | ||
| MRI-BrainDuke (MRI-Brain) | Brain with FLAIR Abnormality Segmentation Masks[10] | ||
| MRI-BrainResection (MRI-Resect.) | EPISURG Segmentation on the Resection CaViTy[66] | ||
| MRI-BrainStrokeLesion (MRI-Stroke.) | ISLES 2022: Stroke Lesion Segmentation[33] | ||
| MRI-BreastDuke (MRI-Breast) | Breast Cancer Screening[71, 52] | ||
| MRI-Heart | Medical Segmentation Decathlon Cardiac[4] | ||
| MRI-Hippocampus (MRI-Hippo.) | Medical Segmentation Decathlon Hippocampus Head and Body[4] | ||
| MRI-Kidney | T2-weighted Kidney MRI[20] | ||
| MRI-Prostate (MRI-Prost.) | Original Multi-Parametric MRI Images of Prostate[50] | ||
| CT-Colon | Medical Segmentation Decathlon Colon[4] | ||
| CT-HepaticVessel (CT-vessel) | Medical Segmentation Decathlon Hepatic Vessel[4] | ||
| CT-Pancreas (CT-Panc.) | Medical Segmentation Decathlon Pancreas[4] | ||
| US-Breast | Breast Ultrasound Images Dataset[2] | ||
| US-Kidney | Kidney Segmentation in Ultrasound Images[77] | ||
| US-Muscle | Transverse Musculoskeletal Ultrasound images[60] | ||
| Xray-Chest | Montgomery County and Shenzhen Chest X-ray Datasets[41] | ||
| Xray-Hip | X-ray Images of the Hip Joints[30] | ||
@@
| MRI-BrainSubcortical | 14 Brain Sub-cortical Structures[76] | ||
| MRI-CHAOS | Segmentation of Liver, Kidneys and Spleen from MRI data[44] | ||
| MRI-Knee | MRNet Knee MRI Examinations[9] | Not used | |
| MRI-ProstateX | ProstateX Dataset of Prostate MR Studies[5] | ||
| MRI-SpineSPIDER | SPIDER - Lumbar Spine Segmentation in MR Images[81] | ||
@@ 2.3 Task-general Prompt-based Learning
When multiple labeled datasets are available, an approach to adapting SAM to the medical imaging field involves additional pre-training, i.e. starting with SAM’s pre-trained weights and continuing the learning process with a compilation of the medical datasets, followed by task-specific supervised training, as is similar in [18, 59]. The procedure is illustrated in Figure 1 setup . The motivation for adopting a two-step training approach, rather than utilizing automated fine-tuning on multiple datasets, lies in the fact that various datasets, despite featuring medical images of the same anatomical area and being captured using the same imaging technique, may concentrate on entirely distinct objectives. For example, two brain MRI datasets might each focus on segmenting different entities, such as the hippocampus in one and tumors in the other. Applying automated segmentation to these datasets could introduce ambiguity, and make the training challenging.
@@
Therefore, we use a prompt-based segmentation setup during task-expansive prompt-based learning, which involves providing prompts to identify the targeted area of interest interactively. This approach addresses the issue of the above-mentioned ambiguity. Furthermore, this choice conceptually aligns with the second phase of SAM’s second-stage training process [45], allowing it to be considered as an extra step of supervised training specifically tailored for medical images within the SAM framework.
@@
To simulate a prompt, we first randomly selected a type of object if the ground truth mask contains multiple objects. Then, for each contiguous region of the selected object, we generate a box or a point prompt with equal probability. The generation of box prompts follows SAM’s strategy: a tight box containing the target box is created first, followed by some random noises added to both locations. The random noises are generated with the standard deviation equal to of the box coordinates and to a maximum of pixels. For point generation, we follow [61] to first compute the distance of each pixel to its closet edge within the object region and sample a pixel with a distance above the median distance. This operation allows us to sample a point prompt around the center of the object. Such a design choice is different from that of the SAM, which samples the first point prompt uniformly and subsequent points from the error region between the previous mask prediction and the ground truth mask. Since our ultimate goal is to learn an automated segmentation algorithm, we do not need the iterative process to refine the predictions based on prompts.
@@
Considering the differences across various medical imaging modalities, alongside the observation that images within the same modality often share similar visual features [46], we emphasize the value of modality-specific dataset knowledge in pre-training, a practice also common in the literature [78]. Specifically, we focus on the supervised pre-training on MRI, which consists of most of the collected datasets.
@@ 2.4 Task-agnostic Self-supervised Learning
When using multiple unlabeled medical image datasets, we can utilize the same idea as in the last section to conduct additional pre-training before applying task-specific supervised training. As the label information is unavailable in this scenario, we adapt the self-supervised learning (SSL) techniques, referred to as task-agnostic self-supervised learning. SSL refers to training a network without any human-collected labels. This is an effective strategy since the number of available datasets can increase dramatically. Popular SSL methods generally fall into two categories: contrastive learning [32, 16, 65] or masked image modeling [31, 90]. In the medical image field, further methods are tailored to fit the characteristics of medical images [78], including dimension discrepancy [89], structure similarity [43], etc.
@@
Among these methods, we select the Masked Autoencoder (MAE) [31] as our self-supervised pre-training strategy for three reasons: (1) the computation of CL is infeasible given the input size of SAM, i.e. ; (2) SAM also utilized an MAE pre-trained vision transformer as the initial model; (3) there is increasing evidence showing the effectiveness of using MAE during medical image segmentation [100, 101, 84]. MAE first divides input images into non-overlapping patches and removes most of them, and its goal is to reconstruct these original inputs given the remaining ones. For the same reason as in task-expansive prompt-based learning, we also conduct training with MRIs only and evaluate all datasets.
3 Dataset
@@
We select 22 datasets from publicly available sources online, including TCIA (The Cancer Imaging Archive), Kaggle, Zendo, Grand Challenge, and Mendeley Data. Figure 3 presents a visual summary of the datasets used in this paper. In total, we collect magnetic resonance images (MRI) datasets, computed tomography (CT) datasets, ultrasound (US) datasets, and X-ray datasets. Our datasets cover images from main human anatomical locations. Examples of all datasets are also included in the same figure, in addition to the image count of each dataset, each anatomical location, and each modality. The details of each dataset are listed in Table 2, including the citation, description, number of images, and number of masks. All datasets employ expert annotated segmentation masks except for the MRI-Knee dataset, which is used exclusively for self-supervised pre-training and thus their numbers of masks are indicated as “not used”.
@@
Since SAM is a 2D-based segmentation network, 3D volumes are first extracted to 2d slices. We convert NifTI, mha, and Dicom volumes by normalizing all pixel values to 0-255 and further upsampling each slice to using nearest-neighbor interpolation to meet SAM’s required input size. The pixel values are repeated three times across the three channels for all grayscale images. We split each dataset into training, validation, and test sets by dividing patients into splits accordingly.
@@
In our study, we utilize this curated medical imaging datasets across various levels of data availability as follows:
@@
4 Experiment
In this section, we conduct various experiments to assess the effectiveness of fine-tuning strategies across multiple datasets and configurations. First and foremost, we analyze the behaviors of these strategies when conducting task-specific supervised training. Next, we examine the effectiveness of applying task-expansive prompt-based learning or task-agnostic self-supervised learning before task-specific supervised training. These experiments are further extended to the few-shot learning and interactive segmentation setting. Throughout the paper, we use the dice similarity score (DSC%) as the evaluation metric.
@@ 4.1.1 Implementation details
When conducting task-specific supervised training, in addition to finding the most effective fine-tuning strategy, we also want to compare the performance against conventional UNet-based algorithms. Therefore, we have chosen three well-known and frequently used automated segmentation models for comparison: UNet [70], nnUNet [40], and SwinUNet [13]:
@@
SwinUNet replaces the encoder of UNet with a swin-transformer [58].
For PEFT with Adapter blocks, we implemented it based on [87], which added adapter blocks to SAM’s attention blocks. In the image encoder, adapter blocks were added into transformer blocks right after the multi-attention head or within the MLP residual block with a scalable parameter [15]. In the default setting, we added Adapter blocks at the first two and last two transformer blocks (layers) in the image encoder. In the mask decoder, Adapters were added after the multi-head attention head. For PEFT with LoRA blocks, we implemented it based on SAMed [96], which inserted LoRA layers into the “query” and “value” projection layers of each transformer block in the image encoder and mask decoder.
In the pre-processing step of our datasets, we removed all data lacking target objects, specifically those images where the corresponding masks only consist of the background, for both training and testing purposes. We also followed the same normalization mean and standard deviation as the original SAM. We utilized dice loss and cross-entropy loss with equal weights. For all experiments across various datasets, we adapted a basic learning rate of 1e-4 with a warmup of iterations and applied AdamW with weight decay for the remaining iterations. The batch size was set to . Histogram equalization and color jitter were the augmentations used during training. The maximum epochs were set to epochs, incorporating an early stopping mechanism that halted training if there was no improvement in the highest validation DSC for consecutive epochs. For settings that require heavy computing resources, such as ViT-H with the vanilla setting, we adopted Distributed Data Parallel (DDP) combined with model parallel to split batches or the model into 2-3 GPUs. The specifics regarding the total number of parameters in the model, the number of parameters subject to updates for each combination, and the computing resources needed are detailed in Table 3.
@@ 4.1.2 Results
Figure 4 shows the performance of various task-specific fine-tuning approaches across each dataset and the performance of various UNet-based algorithms. The detailed results of each architecture-component-method fine-tuning strategy over 17 datasets can be found in the Appendix. Note that we provide the numerical results for all figure-based results in the Appendix. By comparing the red and blue bars, we first observe an almost definite (except in the ViT-B + Vanilla setting) improvement in the average performance when changing from only updating the decoder parameters to that of both the encoder and decoder. This discovery indicates that SAM’s original image encoder, initially trained on a vast collection of natural images, is not well-suited for efficiently extracting features from medical images, aligned with previous studies [61]. Therefore, freezing it and directly adapting the mask decoder for downstream tasks in the medical domain results in sub-optimal performance.
@@
When comparing the performances among Vanilla, Aapter, and LoRA, we observe little difference between the two PEFT methods (Adapter and LoRA), and the rank between Vanilla and PEFT methods depends on the choices of backbones. When using the standard-size backbone (ViT-H or ViT-B), we observe an absolute advantage of using PEFT techniques. For example, when comparing the middle three blue bars in Figure 4, we observe using PEFT with Adapter or LoRA achieves and DSC respectively, while vanilla fine-tuning yields only DSC. We also conduct an additional study to insert the Adapter block after each layer and observe little difference. However, the difference between vanilla fine-tuning and PEFT is significantly reduced when using ViT-T, the small backbone introduced by MobileSAM [94]. We hypothesize the large gap between Vanilla and PEFT methods with a standard-size backbone is caused by the model collapse and knowledge-forgotten issues when fine-tuning all parameters of a large model [23, 26]. To support our hypothesis, we track the change of validation DSC on all datasets, and the examples of each modality are shown in Figure 5. It reveals that ViT-H + En/Decoder + Vanilla would cause a model collapse at an early stage. For instance, when applied to the MRI-BrainDuke dataset, this configuration results in a significant reduction in performance, with a decrease of approximately in DSC.
@@
Next, our results reveal that the choice of encoder network architecture has minimal impact on performance when other choices remain the same. For instance, under the same settings of updating both encoder and decoder with Adapter (EN/Decoder + Adapter), we observe similar DSC scores across different encoders: ViT-H, ViT-B, and ViT-T. This similarity in performance occurs despite ViT-H having a significantly larger encoder with times network parameters compared with ViT-B and times parameters compared with ViT-T, as listed in Table 3. We deduce that while ViT-H excels in foundational settings for multiple segmentation tasks, it may be too parameter-rich for singular medical imaging segmentation tasks, given the typically smaller dataset sizes in the medical imaging field. Such property can also make the training difficult. As shown in Figure 5 top right, it takes epochs to train the ViT-H + En/Decoder + Adapter setting, which is approximately times longer when conducting the same setting with ViT-B, yet the latter gives slightly better performance ( DSC vs. DSC when using ViT-H and ViT-B as the backbone respectively). Given the uniform results across our 17 datasets, we deduce that using PEFT (Adapter or LoRA) to fine-tune SAM within a smaller framework offers more benefits. This strategy strikes an optimal balance between efficiency and performance, especially for medical image segmentation tasks where datasets typically range from a couple hundred to tens of thousands of images.

### v3

@@ 2.2.4 Summary
To sum up, we choose the following combinations to fine-tune SAM’s performance for specific datasets or tasks, as illustrated in Figure 2: (1) Variations in the image encoder’s architecture, including ViT-H, ViT-B, and ViT-T from MobileSAM (Zhang et al., 2023a). (2) Different approaches to the fine-tuning scope of the network, encompassing both the encoder and decoder updates, as well as the decoder only. (3) Diverse parameter updating settings, comprising Vanilla, PEFT with Adapter blocks, and PEFT with LoRA blocks. By integrating these three variables, we present a total of distinct architectures for fine-tuning SAM on task-specific supervised training setup.
@@
| MRI-BrainSclerosis (MRI-Scler.) | Brain Sclerosis with Lesion Segmentation (Muslim et al., 2022) | ||
| MRI-BrainDuke (MRI-Brain) | Brain with FLAIR Abnormality Segmentation Masks (Buda, 2019) | ||
| MRI-BrainResection (MRI-Resect.) | EPISURG Segmentation on the Resection CaViTy (Pérez-García et al., 2020) | ||
| MRI-BrainStrokeLesion (MRI-Stroke.) | ISLES 2022: Stroke Lesion Segmentation (Hernandez Petzsche et al., 2022) | ||
| MRI-BreastDuke (MRI-Breast) | Breast Cancer Screening (Saha et al., 2018; Lew et al., 2024) | ||
| MRI-Heart | Medical Segmentation Decathlon Cardiac (Antonelli et al., 2022) | ||
| MRI-Hippocampus (MRI-Hippo.) | Medical Segmentation Decathlon Hippocampus (Antonelli et al., 2022) | ||
| MRI-Kidney | T2-weighted Kidney MRI (Daniel et al., 2021) | ||
| MRI-Prostate (MRI-Prost.) | Original Multi-Parametric MRI Images of Prostate (Lemaitre et al., 2016) | ||
| CT-Colon | Medical Segmentation Decathlon Colon (Antonelli et al., 2022) | ||
| CT-HepaticVessel (CT-vessel) | Medical Segmentation Decathlon Hepatic Vessel (Antonelli et al., 2022) | ||
| CT-Pancreas (CT-Panc.) | Medical Segmentation Decathlon Pancreas (Antonelli et al., 2022) | ||
| US-Breast | Breast Ultrasound Images Dataset (Al-Dhabyani et al., 2019) | ||
| US-Kidney | Kidney Segmentation in Ultrasound Images (Song et al., 2022) | ||
| US-Muscle | Transverse Musculoskeletal Ultrasound images (Marzola et al., 2021) | ||
| Xray-Chest | Montgomery County and Shenzhen Chest X-ray Dataset (Jaeger et al., 2014) | ||
| Xray-Hip | X-ray Images of the Hip Joints (Gut, 2021) | ||
@@
| MRI-BrainSubcortical | 14 Brain Sub-cortical Structures (Sivaswamy et al., 2021) | ||
| MRI-CHAOS | Segmentation of Liver, Kidneys and Spleen from MRI data (Kavur et al., 2021) | ||
| MRI-Knee | MRNet Knee MRI Examinations (Bien et al., 2018) | Not used | |
| MRI-ProstateX | ProstateX Dataset of Prostate MR Studies (Armato III et al., 2018) | ||
| MRI-SpineSPIDER | SPIDER - Lumbar Spine Segmentation in MRI (van der Graaf et al., 2023) | ||
@@ 2.3 Task-general Prompt-based Learning
When multiple labeled datasets are available, an approach to adapting SAM to the medical imaging field involves additional pre-training, i.e. starting with SAM’s pre-trained weights and continuing the learning process with a compilation of the medical datasets, followed by task-specific supervised training, as is similar in (Cheng et al., 2023; Ma et al., 2024). The procedure is illustrated in Figure 1 setup . The motivation for adopting a two-step training approach, rather than utilizing automated fine-tuning on multiple datasets, lies in the fact that various datasets, despite featuring medical images of the same anatomical area and being captured using the same imaging technique, may concentrate on entirely distinct objectives. For example, two brain MRI datasets might each focus on segmenting different entities, such as the hippocampus in one and tumors in the other. Applying automated segmentation to these datasets could introduce ambiguity and make the training challenging.
@@
Therefore, we use a prompt-based segmentation setup during task-expansive prompt-based learning, which involves providing prompts to identify the targeted area of interest interactively. This approach addresses the issue of the above-mentioned ambiguity. Furthermore, this choice conceptually aligns with the second phase of SAM’s second-stage training process (Kirillov et al., 2023), allowing it to be considered as an extra step of supervised training specifically tailored for medical images within the SAM framework.
@@
To simulate a prompt, we first randomly selected a type of object if the ground truth mask contains multiple objects. Then, for each contiguous region of the selected object, we generate a box or a point prompt with equal probability. The generation of box prompts follows SAM’s strategy: a tight box containing the target box is created first, followed by some random noises added to both locations. The random noises are generated with the standard deviation equal to of the box coordinates and to a maximum of pixels. For point generation, we follow (Mazurowski et al., 2023) to first compute the distance of each pixel to its closest edge within the object region and sample a pixel with a distance above the median distance. This operation allows us to sample a point prompt around the center of the object. Such a design choice differs from that of the SAM, which samples the first point prompt uniformly and subsequent points from the error region between the previous mask prediction and the ground-truth mask. Since our ultimate goal is to learn an automated segmentation algorithm, we don’t need an iterative process to refine predictions based on prompts.
@@
Considering the differences across various medical imaging modalities, alongside the observation that images within the same modality often share similar visual features (Konz and Mazurowski, 2024), we emphasize the value of modality-specific dataset knowledge in pre-training, a practice also common in the literature (Tang et al., 2022). Specifically, we focus on supervised pre-training on MRI, which comprises most of the collected datasets.
@@ 2.4 Task-agnostic Self-supervised Learning
When using multiple unlabeled medical image datasets, we can utilize the same idea as in the last section to conduct additional pre-training before applying task-specific supervised training. As the label information is unavailable in this scenario, we adapt the self-supervised learning (SSL) techniques, referred to as task-agnostic self-supervised learning. SSL refers to training a network without any human-collected labels. This is an effective strategy since the number of available datasets can increase dramatically. Popular SSL methods generally fall into two categories: contrastive learning (He et al., 2020; Chen et al., 2020; Park et al., 2020) or masked image modeling (He et al., 2022; Xie et al., 2022b). In the medical image field, further methods are tailored to fit the characteristics of medical images (Tang et al., 2022), including dimension discrepancy (Xie et al., 2022a), structure similarity (Jiang et al., 2023), etc.
@@
Among these methods, we select the Masked Autoencoder (MAE) (He et al., 2022) as our self-supervised pre-training strategy for three reasons: (1) The computation of CL is infeasible given the input size of SAM, i.e. ; (2) SAM also utilized an MAE pre-trained vision transformer as the initial model; (3) There is increasing evidence showing the effectiveness of using MAE during medical image segmentation (Zhou et al., 2023; Zhuang et al., 2023; Wang et al., 2024). MAE first divides the input images into non-overlapping patches and removes most of them. Its goal is to reconstruct the original inputs given the remaining ones. For the same reason as in task-expansive prompt-based learning, we also conduct training with MRIs only and evaluate all datasets.
3 Datasets
@@
We select 22 datasets from publicly available sources online, including TCIA (The Cancer Imaging Archive), Kaggle, Zendo, Grand Challenge, and Mendeley Data. Figure 3 presents a visual summary of the datasets used in this paper. In total, we collect magnetic resonance images (MRI) datasets, computed tomography (CT) datasets, ultrasound (US) datasets, and X-ray datasets. Our datasets cover images from main human anatomical locations. Examples of all datasets are also included in the same figure, in addition to the image count of each dataset, each anatomical location, and each modality. The details of each dataset are listed in Table 2, including the citation, description, number of images, and number of masks. All datasets use expert-annotated segmentation masks, except for the MRI-Knee dataset, which is used exclusively for self-supervised pre-training. Therefore, the number of masks for this dataset is indicated as “not used”.
@@
Since SAM is a 2D-based segmentation network, 3D volumes are first extracted to 2D slices. We convert NIfTI, MHA, and DICOM volumes by normalizing all pixel values to 0-255 and further upsampling each slice to using nearest-neighbor interpolation to meet SAM’s required input size. The pixel values are repeated three times across the three channels for all grayscale images. We split each dataset into training, validation, and test sets by dividing patients into splits, accordingly.
@@
In our study, we utilize these curated medical imaging datasets across various levels of data availability as follows:
@@
4 Experiments
In this section, we conduct various experiments to assess the effectiveness of fine-tuning strategies across multiple datasets and configurations. First and foremost, we analyze the behaviors of these strategies when conducting task-specific supervised training. Next, we examine the effectiveness of applying task-expansive prompt-based learning or task-agnostic self-supervised learning before task-specific supervised training. These experiments are further extended to the few-shot learning and interactive segmentation setting. Throughout the paper, we use the Dice similarity score (DSC%) and normalized surface distance (NSD%) as the evaluation metric.
@@ 4.1.1 Implementation details
When conducting task-specific supervised training, in addition to finding the most effective fine-tuning strategy, we also want to compare the performance against conventional UNet-based algorithms. Therefore, we have chosen three well-known and frequently used automated segmentation models for comparison: UNet (Ronneberger et al., 2015), nnUNet (Isensee et al., 2021), and SwinUNet (Cao et al., 2022):
@@
SwinUNet replaces the encoder of UNet with a swin-transformer (Liu et al., 2021).
For PEFT with Adapter blocks, we implemented it based on (Wu et al., 2023b), which added adapter blocks to SAM’s attention blocks. In the image encoder, adapter blocks were added into transformer blocks right after the multi-attention head or within the MLP residual block with a scalable parameter (Chen et al., 2022). In the default setting, we added Adapter blocks at the first two and last two transformer blocks (layers) in the image encoder. In the mask decoder, Adapters were added after the multi-head attention. For PEFT with LoRA blocks, we implemented it based on SAMed (Zhang and Liu, 2023), which inserted LoRA layers into the “query” and “value” projection layers of each transformer block in the image encoder and mask decoder.
In the pre-processing step for our datasets, we removed all data lacking target objects, specifically images where the corresponding masks only consisted of the background, for both training and testing purposes. We also followed the same normalization mean and standard deviation as the original SAM. We utilized dice loss and cross-entropy loss with equal weights. For all experiments across various datasets, we adopted a basic learning rate of 1e-4 with a warmup of iterations and applied AdamW with weight decay for the remaining iterations. The batch size was set to . Histogram equalization and color jitter were used as augmentations during training. The maximum epochs were set to epochs, incorporating an early stopping mechanism that halted training if there was no improvement in the highest validation DSC for consecutive epochs. For settings that require heavy computing resources, such as ViT-H with the vanilla setting, we used Distributed Data Parallel (DDP) combined with model parallelism to split batches or the model across 2-3 GPUs. The specifics regarding the total number of parameters in the model, the number of parameters subject to updates for each combination, and the computing resources needed are detailed in Table 3.
@@ 4.1.2 Results
Figure 4 shows the performance of various task-specific fine-tuning approaches across each dataset and the performance of various UNet-based algorithms. The detailed results of each architecture-component-method fine-tuning strategy over 17 datasets can be found in the Appendix. Note that we provide the numerical results for all figure-based results in the Appendix. By comparing the red and blue bars, we first observe an almost definite improvement (except in the ViT-B + Vanilla setting) in average performance when updating both the encoder and decoder parameters. This discovery indicates that SAM’s original image encoder, initially trained on a vast collection of natural images, is not well-suited for efficiently extracting features from medical images, aligned with previous studies (Mazurowski et al., 2023). Therefore, freezing it and directly adapting the mask decoder for downstream tasks in the medical domain yields suboptimal performance.
@@
When comparing the performances of Vanilla, Adapter, and LoRA, we observe little difference between the two PEFT methods (Adapter and LoRA). The rank between Vanilla and the PEFT methods depends on the choices of backbones. When using the standard-size backbone (ViT-H or ViT-B), we observe an absolute advantage of using PEFT techniques. For example, when comparing the middle three blue bars in Figure 4, we observe that using PEFT with Adapter or LoRA achieves and DSC, respectively, while vanilla fine-tuning yields only DSC. When comparing Adapter and LoRA, we observe that LoRA performs better when applied to a model with a larger size, i.e., ViT-H. This can be caused by the superior parameter efficiency of LoRA (Hu et al., 2021). We also conducted an additional study to insert the adapter block after each layer and observed little difference. However, the difference between vanilla fine-tuning and PEFT is significantly reduced when using ViT-T, the small backbone introduced by MobileSAM (Zhang et al., 2023a). We hypothesize that the large gap between Vanilla and PEFT methods with a standard-size backbone is caused by the model collapse and knowledge-forgotten issues when fine-tuning all parameters of a large model (Dutt et al., 2023; Fu et al., 2023). To support our hypothesis, we track the change in validation DSC on all datasets. Examples of each modality are shown in Figure 5. It reveals that ViT-H + En/Decoder + Vanilla would cause a model collapse at an early stage. For instance, when applied to the MRI-BrainDuke dataset, this configuration results in a significant performance decrease, with a drop of approximately in DSC.
@@
Next, our results show that the choice of encoder network architecture has a minimal impact on performance when all other decisions remain the same. For instance, under the same settings of updating both encoder and decoder with Adapter (EN/Decoder + Adapter), we observe similar DSC scores across different encoders: ViT-H, ViT-B, and ViT-T. This similarity in performance occurs despite ViT-H having a significantly larger encoder with times network parameters compared with ViT-B and times parameters compared with ViT-T, as listed in Table 3. We deduce that while ViT-H excels in foundational settings for multiple segmentation tasks, it may be too parameter-rich for singular medical imaging segmentation tasks, given the typically smaller dataset sizes in the medical imaging field. Such property can also make the training difficult. As shown in Figure 5 top right, it takes epochs to train the ViT-H + En/Decoder + Adapter setting, which is approximately times longer when conducting the same setting with ViT-B, yet the latter gives slightly better performance ( DSC vs. DSC when using ViT-H and ViT-B as the backbone respectively). Given the uniform results across our 17 datasets, we deduce that using PEFT (Adapter or LoRA) to fine-tune SAM within a smaller framework offers more benefits. This strategy strikes an optimal balance between efficiency and performance, especially for medical image segmentation tasks where datasets typically range from a couple of hundred to tens of thousands of images.